OpenStudy (anonymous):

Translate the situation into an equation... Alice is twice as old as Lisa and Brandon is four years younger than Lisa. If their average age age is 32, how old is Alice.. Please show steps

OpenStudy (anonymous):

Ready to work on this together?

OpenStudy (anonymous):

Let me know whenever you're here. We'll get the answer together so that you'll always know how to solve these!

OpenStudy (anonymous):

im here

OpenStudy (anonymous):

Awesome. Ok, first things first: who are we using as our reference point? In other words, who are we using to compare to the other people? If that question doesn't make sense, I can reword it.

OpenStudy (anonymous):

i think Alice

OpenStudy (anonymous):

We're actually using Lisa. Notice how we say that alice is twice as old as Lisa, and that brandon is four years younger than Lisa. If we know how old Lisa is, then we'll know how old the other two are. Does that make sense?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Okie dokie. So what we'll do is say that Lisa = x, since she's the 'unknown age that we will use to compare everyone to her"

OpenStudy (anonymous):

Remember: lisa is now called "x".

OpenStudy (anonymous):

Alice is TWO times older than Lisa, I mean.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Alright, so show me what Alice's age would be represented by.

OpenStudy (anonymous):

2x

OpenStudy (anonymous):

Perfect! So we know that lisa = x and alice is 2x Now all we have left is Brandon. He is four years younger than Lisa. How do we show this?

OpenStudy (anonymous):

-4x

OpenStudy (anonymous):

That would actually mean "minus four times". Note that we are just saying he's four years younger, not four times younger.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

In other words, "If I have 10 apples, and Trust26 has 4 less apples than me, how do we show this in an equation?"

OpenStudy (anonymous):

10 -4

OpenStudy (anonymous):

Perfect. So in Brandon's case, we have x-4, right?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Alright! So here's where it gets a little tricky. Imagine I told you to find the average of 10, 4 and 1. What would you do?

OpenStudy (anonymous):

Solve the following simultaneous equations for alice, brandon and lisa: \[\left\{\text{alice}\text{=}2\text{ lisa} ,\text{brandon}=\text{lisa}-4,\frac{\text{alice}+\text{brandon}+\text{lisa}}{3}\text{=}32\right\} \]

OpenStudy (anonymous):

^ I think that might be a little tricky for an algebra 1 student.

OpenStudy (anonymous):

O am not for sure Uber

OpenStudy (anonymous):

You would add all three first, right?

OpenStudy (anonymous):

10 + 4 + 1 = 15

OpenStudy (anonymous):

Now, what is the second step?

OpenStudy (anonymous):

ok imma lost where did you get 10+4+1 from

OpenStudy (anonymous):

Oh, that's just an example I'm making to show you what you will do later.

OpenStudy (anonymous):

I'm checking to make sure you know how to average stuff before we go back to the problem itself. The 10,4,1 are just made up numbers I'm going to use to show you the steps for making an average.

OpenStudy (anonymous):

ok. i got you. go ahead

OpenStudy (anonymous):

Alright. So do you know what you do after you get 15?

OpenStudy (anonymous):

divide

OpenStudy (anonymous):

Excellent. By what number?

OpenStudy (anonymous):

\[\{\text{alice}\to 50,\text{brandon}\to 21,\text{lisa}\to 25\} \]

OpenStudy (anonymous):

3

OpenStudy (anonymous):

Exactly. So we'd have (10+4+1)/3 = 5, where 5 is the average. Now... we're going to do the same thing for your original question.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

( )/[ ] = {} Our formula will look like that, kinda. Do you see how the blanks match the "(10+4+1)/[3] = {5}" question?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Alice is twice as old as Lisa and Brandon is four years younger than Lisa. If their average age age is 32, how old is Alice

OpenStudy (anonymous):

Ok, so what do we put in the ( ) ?

OpenStudy (anonymous):

Hint: the ( ) in our first example was 10 + 1 + 4

OpenStudy (anonymous):

we put their ages in the ( )

OpenStudy (anonymous):

Exactly! So here's what we had: (Alice + Lisa + Brandon) (x + 2x + x-4) So far so good?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Okie dokie. First thing we should do is simplify the junk in there. What would we get after combining like terms? (x + 2x + x-4)

OpenStudy (anonymous):

4x-4 than divide 4x/4 and 4/4

OpenStudy (anonymous):

It's better if we just leave it at 4x-4 to be honest. We shouldn't divide anything until we have something in the form of 4x-4 = "something". So, we have (4x-4)/[] = {}

OpenStudy (anonymous):

How many people do we have?

OpenStudy (anonymous):

3

OpenStudy (anonymous):

Excellent. So we have (4x-4)/3 = {} What goes in the {}?

OpenStudy (anonymous):

Alice is twice as old as Lisa and Brandon is four years younger than Lisa. If their average age age is 32, how old is Alice..

OpenStudy (anonymous):

50

OpenStudy (anonymous):

We'd actually put in 32, as that was the answer to our average.

OpenStudy (anonymous):

(4x-4)/3 = 32 Make sense?

OpenStudy (anonymous):

ok ok yes i see

OpenStudy (anonymous):

Awesome. So here's the next part: we want all our X's to be on one side, and our other numbers on the other. Do you know how we can get rid of the /3 first?

OpenStudy (anonymous):

Here's a hint: If I had 10/2 = 5, how do I move the 2 to the other side while still making it a true statement?

OpenStudy (anonymous):

(That's another random example not directly related to our equation, by the way)

OpenStudy (anonymous):

add whats in the ( ) first

OpenStudy (anonymous):

Well, we can't get any deeper than (4x+4)

OpenStudy (anonymous):

Try that other example first. 10/2 = 5 How do we move the 2 to the right and still let it be true? As in... 10 = 2 [] 5

OpenStudy (anonymous):

What mathematical symbol goes in that []

OpenStudy (anonymous):

divide

OpenStudy (anonymous):

Careful.

OpenStudy (anonymous):

10 = 2/5 is not true

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

So 5 [what] 2 = 10? +? -? /? *?

OpenStudy (anonymous):

o times

OpenStudy (anonymous):

excellent.

OpenStudy (anonymous):

So remember: If you ever have something like a/b = c, and you need to get a by itself, just move the b to the right and times it. So in this case, a = b * c. Makes sense?

OpenStudy (anonymous):

yes it does

OpenStudy (anonymous):

Awesome. (4x-4)/3 = 32 How do we make the 3 go to the other side?

OpenStudy (anonymous):

(Don't worry. We're *almost* done.)

OpenStudy (anonymous):

just move it

OpenStudy (anonymous):

Yup. And what else?

OpenStudy (anonymous):

and times it

OpenStudy (anonymous):

Exactly. So what do we get? (4x-4)/3 = 32

OpenStudy (anonymous):

(4x-4)= 32*3

OpenStudy (anonymous):

Excellent. So 4x-4 = 96 4x = 100 x = 25 Did these steps make sense?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Translate the situation into an equation... Alice is twice as old as Lisa and Brandon is four years younger than Lisa. If their average age age is 32, how old is Alice.. Please show steps

OpenStudy (anonymous):

So, is 25 our final answer?

OpenStudy (anonymous):

ok i ok see no cause you have to time it by two

OpenStudy (anonymous):

ok i ok see no cause you have to time it by two

OpenStudy (anonymous):

Perfect! So, 50. :) Good work.

OpenStudy (anonymous):

o thank yoou for the help you deserve a medal to and i am going to give it to you. beacsue you actually took time out to teac me

OpenStudy (anonymous):

Yay :D

OpenStudy (anonymous):

And thank YOU for trying to solve unlike some folks that just say "No. It's too hard. I give up." :D